OpenStudy (anonymous):

What is 3/5 written as a decimal? A.0.3 B.0.6 C.0.12 D.1.67

OpenStudy (alexandervonhumboldt2):

3/5*2=6/10=0.6

OpenStudy (faithhouse):

B.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Its has to be bigger than half, so not .3,

OpenStudy (anonymous):

Thanks @FaithHouse and @AlexandervonHumboldt2

OpenStudy (faithhouse):

the simple way to do it is taking a calculator and divided 3 by 5 = 0.6

OpenStudy (faithhouse):

Much faster

OpenStudy (anonymous):

Brb

OpenStudy (faithhouse):

k

OpenStudy (anonymous):

thanks for your help got to close the question now